20k of water is evaporated from a vat that contains 200kg of an 11% brine solution (11% salt,89% water). 5 kg of salt is then added to the brine. The brine solution that results from this process will contain what percentage of salt?
Answer by richwmiller(17219)

200*.11=(200-20)*a,a = 0.122222
200*.89-20=180*b, b=0.877778, a=0.122222
180*.122222+5=185*c,c = 0.145946
200*.11+5=c(200-20+5),c = 0.145946
final 14.5946 %